the molor solubility for Pb3(po4)2 is 7.7 x 10^-10 moles L ^-1. calculate kap
Answer #1

Answer

Ksp = 2.92 × 10-44

Explanation

Solubility equilibrium of Pb3(PO4)2 is

Pb3(PO4)2(s) <-------> 3Pb2+(aq) + 2PO43-(aq)

Ksp = [Pb2+]3[PO43-]2

at saturated solution

[Pb2+] = 3 × 7.7 ×10-10M = 2.31 × 10-9 M

[PO43-] = 2 × 7.7 ×10-10M = 1.54 ×10-9M

Ksp = ( 2.31 ×10-9)3 ( 1.54 ×10-9)2

Ksp = 2.92 × 10-44